Construct a 90% confidence interval for the true proportion of all employees of Big Star Ltd who believe that they deserve a salary increase, given that 80 out of a sample of 130 employees said that they believe that they deserve a salary increase.


1
Expert's answer
2021-05-26T03:47:51-0400

We have given that,


"n = 130"


"p = \dfrac{80}{130} = 0.615"


We know,


"CI = p \pm z \sqrt{\dfrac{p(1-p)}{n}}"


"CI = 0.615 \pm 1.645\sqrt{\dfrac{0.615 \times 0.385}{130}}"


"CI = 0.615 \pm 0.0070"